A compound containing only C, H, and O, was extracted from the bark of the sassafras tree. The combustion of 44.3 mg produced 120 mg of CO2 and 24.6 mg of H2O. The molar mass of the compound was 162 g/mol. Determine its empirical and molecular formulas.
1
Expert's answer
2020-02-03T07:45:52-0500

n(CxHyOz)=44.3/162=0.273mmoles

n(CO2)=120/44=2.73mmoles

n(H2O)=24.6/18=1.37mmoles

x=n(CO2)/n(CxHyOz)=10

y=2*n(H2O)/n(CxHyOz)=10

z=(162-12*10-1*10)/16=2.